# Formal Definition and Implementation of Reproducibility Tenets for Computational Workflows

### v1

@@ 4.1.1 Rerun
A workflow reruns another if they execute the same logical workflow; that is, their logical components and dependencies match.
Rerunning has been suggested to show that used tools are robust; rerunning a computational workflow is analogous to using an identical algorithm independent of its implementation [39]. Asserting consistency only on the logical components in a workflow allows physical component changes. This definition, in turn, allows changes in computing scale, parallelism patterns, hardware, data-encoding, or software implementation while still constituting a rerun. If all physical details are kept identical, a successful rerun builds trust in the underlying tooling used to deploy a workflow, and a failed rerun demonstrates system failure. Rerunning workflows is useful when scaling up from a laptop to a cluster, for instance, or for workflows that, by design, deal with wildly varying data inputs, such as an ingest pipeline.
@@ 4.1.2 Repeat
A workflow repeats another if they execute the same logical workflow and a principally identical physical workflow; their logical components and dependencies and physical tasks match. The specificity of a ‘principally identical’ physical workflow allows for a successful recovery from failed physical tasks (due to hardware or operating failure).
Repetition is a relatively well-defined concept, and we generally follow intuition. Repetition builds statistical power behind results [39]. By analogy to laboratory work, repeated scientific computational workflows should match all components (workflow specification, data input, and runtime environment) [38]. Our definition mostly follows existing interpretations but provides a concrete definition. Repeating a scientific computational workflow asserts the reuse of the logical and physical workflows, which, if successful, allows the bundling of data artifacts in concert. For instance, survey analysis runs on many patches of sky to create a full catalog, repeating the same workflow. Repetition has also been known as methods reproducibility [40]. Importantly, we leave the amount of information clarifying physical tasks as an environment and problem-dependent decision but insist that hardware details are not pertinent for repetition.
@@ 4.1.3 Recompute
A workflow recomputes another if they execute the same physical workflow; that is, their physical tasks and dependencies match precisely.
Recomputation is a stricter interpretation of computational workflow repetition that requires identical physical tasks and hardware and is not fault-tolerant. Workflow provenance, taken to an albeit extreme conclusion, motivates this definition; the effort required to recompute a given workflow is immense and is likely impossible without total hardware control. Some forms of repetition [11] are similar as they call principally for removing non-determinism. In distributed computational workflows, the distribution itself can cause remarkably complex numerical non-determinism that is painstaking to correct [41], so recomputation is a lofty endeavor. Recomputation can be useful, however, to demonstrate that a new computing facility is stable and to isolate platform-dependent difficulties or errors.
@@ 4.1.4 Reproduce
A workflow reproduces another if their scientific information match.
Scientific reproduction is a widely discussed topic, and our definition does not conflict with existing literature. Our definition of reproduction captures the net data flow a computational data-processing workflow consumes and produces. Generally, reproducibility proper refers to reproducing experimental results [42], an extremely difficult task in lab-based science [43]. Some define reproducible codes to be trivially executable and to produce published results [11]. Our definition focuses solely on data in line with other interpretations [39, 5] as opposed to more conceptual definitions based on conclusions [38]. The many interpretations of reproducibility present a challenge with many valid but incompatible definitions [6].
Focussing on data artifacts specifically provides an actionable approach to testing scientific reproducibility. The formal reproduction of other teams’ work and the reproduction of historical but impactful results are healthy for the wider academic community. Our approach is also compatible with reproducibility approaches involving constrained computing environments [44] since we place no constraint on logical or physical task components.
@@ 4.1.5 Replicate
Replicability in science is a substantially debated concept with many interpretations [42, 6]. We incorporate replication into our framework as combinations of our previous tenets, creating three different but related standards that all constrain workflow tasks and data artifacts.
A workflow scientifically replicates another if they rerun and reproduce each other.
@@
Our definition asserts that a principally identical computational workflow produces the same terminal data artifacts, in effect verifying that a logical workflow can produce a given set of results. Our interpretations are reminiscent of what is termed conceptual interpretation [42] and of replicable code [11]. Related but alternate views [38] differ in that we do not focus on the conclusions a workflow implies. For example, scientific replication is useful when updating library versions for a particular workflow, using different implementations for the same tasks or moving to a new computing facility.
A workflow computationally replicates another if they recompute and reproduce each other and all data artifacts match.
@@
Our definition asserts that a precisely identical computational workflow produces the same data artifacts at every stage, which is the strictest interpretation of replication similar to the concept of direct replication [42], and is closely related to workflow provenance discussed widely in computational workflow literature [34]. The information required to replicate a workflow execution computationally is immense. It demonstrates that a specific workflow implementation on specific hardware yields a given set of results, building confidence in workflow implementation but not necessarily the underlying science; it would be challenging for an independent team to totally replicate a complex workflow. Recomputation can confirm anomalous or unexpected results before further verification, for instance.
@@
A workflow totally replicates another if they repeat and reproduce each other and all data artifacts match.
@@
Our definition asserts that a practically identical computational workflow produces the same data artifacts, including interim data, and are identical down to software, but not hardware, details. Replication is useful to certify a workflow as a ‘gold-standard’ [39] and is similar to the concept of results reproducibility [40]. For example, in radio astronomy, an integral processing task requires separating unwanted radio frequency interference from astronomical observations, known as RFI flagging. Not only is the correctness of this task important, but the speed of execution also matters significantly, as this task occurs before any science processing can commence. There exist a multitude of possible methods spanning traditional algorithms and machine learning and a performance claim around a new RFI flagging method could be strengthened by replicating the result independently on different hardware. A total replication permits independent verification of results by a third party and balances scientific and computational replication.
5 Methods
@@
Having provided the theoretical background in the previous section, in this section, we outline our practical implementation, which allows us to verify each of the reproducibility tenants. This includes the specific details behind our workflow signature mechanism, the information captured to build these signatures in DALiuGE, and the details of our demonstrative lowpass filter workflow.
5.1 Blockchain-Based Workflow Signature
@@
Building a dynamic workflow signature is central to our reproducibility tests. Moreover, at SKA scale, these signatures need very efficient construction and comparison routines. To this end, we take inspiration from blockchain technologies, borrowing the Merkle tree cryptographic data structure [45]. Information about each workflow model component, dictated by our tenet definitions, is stored in a Merkle tree, the root of which provides a signature for that component. Inserting these signatures into a blockchain mirroring their connectivity in the original workflow creates a block directed acyclic graph (BlockDAG). The root of a Merkle tree constructed from the leaves of this BlockDAG is a signature for the entire workflow. Similar to how blockchains are cryptographically secure, subtle changes to component provenance information result in drastic changes to the final signature, but structuring this signature makes it efficient to identify where these changes occur. Given two workflow BlockDAGs, one can walk through a topological sort of the nodes to find the first differing signature.
@@
We use Kahn’s algorithm [46] to traverse a workflow description in a topological order taking $\mathcal{O}(V+E)$ time where $V$ is the number of components (vertices) in the workflow and $E$ is the number of edges between all components. Construction of each component Merkle tree takes $\mathcal{O}(|V|\,log\,|V|)$ time where $\|V\|$ is the volume of provenance information the component stores. Inserting a component into the BlockDAG takes $\mathcal{O}(d\,log\,d)$ time, where $d$ is the degree of that component. Constructing the entire BlockDAG takes $\mathcal{O}(V(D\,log,D)+E)$ time, where $D$ is the average vertex degree. Construction of the final signature takes $\mathcal{O}(l\,log\,l)$ time, where $l$ is the number of leaves in the workflow graph. When amortized against the time taken to execute a workflow in the first place ($\mathcal{O}(V+E)$), signature construction takes $\mathcal{O}(log\,l)$ time.
@@ 5.2 Tenet Data
For each of our tenets, we elaborate on what information specifically is important in the case of DALiuGE workflows. Table 3 summarizes the required information for each component category at each workflow abstraction and each tenet. The fields we select are ultimately an interpretation of our tenet definitions and, as such, are open to adjustment, especially when considering alternate workflow management systems or native software.
@@
Rerunning two workflow executions requires matching their logical workflow components. Unsurprisingly, most information captured in this tenet is at the logical layers, only tracking the types and ultimate execution status of each physical component.
@@
In addition to the information stored for asserting reruns, repetition stores additional information about component fields (adjustable parameters).
@@
Recomputation captures significantly more details of component field values and execution environment. In addition to the component details, we store the logical partitioning of components to machine resources and an execution trace for each processing component.
@@
Reproduction significantly differs, primarily focusing on data components. We store information to assert that data components are configured identically and store the same data summary. How components generate their data summary is a subtlety worth noting; in our case, this is a hash of the data contents but could be implemented as calculating statistical properties or other domain-specific characteristics.
@@
Replication information is straightforwardly a combination of the Rerun, Recompute, or Repetition information (respectively) and reproduction fields.
Table 3: Tenet data for each DALiuGE component category, workflow abstraction, and reproducibility tenet. In and Out Ports are the DALiuGE abstraction for a dataflow between components. Node and Island define the logical partitioning of components to compute resources to a specific machine and groups of machines, respectively. Status refers to execution status, trace refers to the stack trace of that component, and not all fields are present in all components, as is the case for Filenames in application components, for instance.
Abstraction Tenet Rerun Repeat Recompute Reproduce Replicate-Sci Replicate-Comp Replicate-Total
LGT Field Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control Application Data Control
Type Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y Y
InPorts Y Y Y Y Y Y Y Y Y - - - Y Y Y Y Y Y Y Y Y
OutPorts Y Y Y Y Y Y Y Y Y - - - Y Y Y Y Y Y Y Y Y
LG Num-CPUs - - - Y - - Y - - - - - - - - Y - - Y - -
Field-Keys - - - Y Y Y Y Y Y - - - - - - Y Y Y Y Y Y
Field-Values - - - - - - Y Y Y - - - - - - Y Y Y - - -
Data-Volume - - - - Y - - Y - - - - - - - - Y - - Y -
Filenames - - - - - - - Y - - - - - - - - Y - - - -
PGS Type Y Y Y Y Y Y Y Y Y - Y - Y Y Y Y Y Y Y Y Y
Name Y - - Y - - Y - - - - - Y - - Y - - Y - -
Storage-Type - Y - - Y - - Y - - Y - - Y - - Y - - Y -
Rank - - - - - - Y Y Y - - - - - - Y Y Y - - -
PGT Node - - - - - - Y Y Y - - - - - - Y Y Y - - -
Island - - - - - - Y Y Y - - - - - - Y Y Y - - -
PG Node-IP-Address - - - - - - Y Y Y - - - - - - Y Y Y - - -
Island-IP-Address - - - - - - Y Y Y - - - - - - Y Y Y - - -
RG Status Y Y - Y Y - Y Y - - Y - Y Y - Y Y - Y Y -
Trace - - - - - - Y - - - - - - - - Y - - - - -
Data-Summary - - - - - - - - - - Y - - Y - - Y - - Y -

### v2

@@ 4.1.1 Rerun
A workflow Reruns another if they execute the same logical workflow; that is, their logical components and dependencies match. Figure 2 shows a version of the generic workflow model highlighting the elements required to be kept identical for workflow Reruns, in this case, logical components without instance-specific fields. We intend that Reruns keep methods of computation identical in principle but open to multiple implementations.
Figure 2: A UML workflow model with required identical components for workflow Reruns highlighted in blue.
Rerunning has been suggested to show that used tools are robust; Rerunning a computational workflow is analogous to using an identical algorithm independent of its implementation [44]. Asserting consistency only on the logical components in a workflow allows physical component changes. This definition, in turn, allows changes in computing scale, parallelism patterns, hardware, data-encoding, or software implementation while still constituting a Rerun. If all physical details are kept identical, a successful Rerun builds trust in the underlying tooling used to deploy a workflow, and a failed Rerun demonstrates system failure. For some concrete examples, Rerunning workflows is useful when scaling up from a laptop to a cluster, for instance, or for workflows that, by design, deal with wildly varying data inputs, such as astronomy observatory ingest pipelines.
@@ 4.1.2 Repeat
A workflow Repeats another if they execute the same logical workflow and a principally identical physical workflow; their logical components, dependencies, and physical tasks match. Figure 3 shows which elements in the workflow model must be identical for two workflow executions to be considered repetitions. The specificity of a ‘principally identical’ physical workflow allows for a successful recovery from failed physical tasks (due to hardware or operating failure). Repetitions are significantly more stringent than Reruns but leave the most specific details of physical tasks, such as the exact machine details these tasks run on and the content of the workflow’s data artifacts.
Figure 3: A UML workflow model with required identical components for workflow repetitions highlighted in blue.
Repetition is a relatively well-defined concept, and we generally follow intuition. Repetition builds statistical power behind results [44]. By analogy to laboratory work, Repeated scientific computational workflows should match all components (workflow specification, data input, and runtime environment) [51]. This definition mostly follows existing interpretations but provides a concrete definition in which all logical and physical components match, excluding the runtime parameters of physical tasks. Repeating a scientific computational workflow asserts the reuse of the logical and physical workflows, which, if successful, allows the bundling of data artifacts in concert. For example, astronomical survey analysis runs on many patches of sky to create a full catalog, Repeating the same workflow but with different data inputs and outputs. Repetition has also been known as methods reproducibility [52]. Importantly, we leave the amount of information clarifying physical tasks as an environment and problem-dependent decision but insist that hardware details are not pertinent for repetition.
@@ 4.1.3 Recompute
A workflow Recomputes another if they execute the same physical workflow; that is, their physical tasks and dependencies match precisely. Figure 4 shows the generic workflow model with all elements required for workflow Recomputation highlighted in blue. The only element not required to be identical is the content of data artifacts.
Figure 4: A UML workflow model with required identical components for workflow Recomputations highlighted in blue.
Recomputation is a stricter interpretation of computational workflow repetition that requires identical physical tasks and hardware and is not fault-tolerant. Workflow provenance, taken to an albeit extreme conclusion, motivates this definition; the effort required to Recompute a given workflow is immense and is likely impossible without total hardware control. Some forms of repetition [11] are similar as they call principally for removing non-determinism. In distributed computational workflows, the distribution itself can cause remarkably complex numerical non-determinism that is painstaking to correct [53], so Recomputation is a lofty endeavor. However, Recomputation can be useful to demonstrate that a new computing facility is stable and isolate platform-dependent difficulties or errors, for example.
@@ 4.1.4 Reproduce
A workflow Reproduces another if their scientific information match. Figure 5 depicts the generic workflow model with components required for workflow reproduction highlighted in blue. Only the logical description of data artifacts and the eventual data a workflow produces is required. The intention is to allow for entire methodologies to change, so long as the outputs of a workflow are maintained, it is a Reproduction, reaching the same end by any means.
Figure 5: A UML workflow model with required identical components for workflow Reproductions highlighted in blue.
Scientific Reproduction is a widely discussed topic; our definition does not conflict with existing literature. This definition of reproduction captures the net data flow a computational data-processing workflow consumes and produces. Generally, reproducibility proper refers to reproducing experimental results [54], an extremely difficult task in lab-based science [55]. Some define reproducible codes to be trivially executable and to produce published results [11]. Our definition focuses solely on data in line with other interpretations [44, 5] as opposed to more conceptual definitions based on conclusions [51]. The many interpretations of reproducibility present a challenge with many valid but incompatible definitions [6].
Focusing on data artifacts provides an actionable approach to testing scientific reproducibility. The formal reproduction of other teams’ work and the reproduction of historical but impactful results are healthy for the wider academic community. Moreover, reproduction alone permits absolute flexibility in methodology; demonstrating reproduction would be useful when establishing a radically different method to perform a known task, as is often the case when introducing machine learning methods, for example. A tenet-testing-based approach is also compatible with reproducibility approaches involving constrained computing environments [56] since we place no constraint on logical or physical task components.
@@ 4.1.5 Replicate
Replicability in science is a substantially debated concept with many interpretations [54, 6]. We incorporate replication into our framework as combinations of previous tenets, creating three different but related standards that all constrain workflow tasks and data artifacts.
A workflow Scientifically Replicates another if they Rerun and Reproduce each other. Figure 6 shows the generic workflow model with components required for workflow scientific replication highlighted in blue. Logical components and the contents of data artifacts need to be identical. Scientific replication demonstrates that a particular method, represented as a logical workflow, produces a specific set of results.
Figure 6: A UML workflow model with required identical components for workflow scientific replications highlighted in blue.
@@
This definition asserts that a principally identical computational workflow produces the same terminal data artifacts, in effect verifying that a logical workflow can produce a given set of results. Our interpretations are reminiscent of what is termed conceptual interpretation [54] and of replicable code [11]. Related but alternate views [51] differ in that we do not focus on the conclusions a workflow implies. For example, scientific replication is useful when updating library versions for a particular workflow, using different implementations for the same tasks, or moving to a new computing facility.
A workflow Computationally Replicates another if they Recompute and Reproduce each other and all data artifacts match. Figure 7 contains the generic workflow model with components required for computational replication highlighted in blue. It should be abundantly clear that computational replication is the most stringent possible form of replication permitted, as all elements must be identical.
Figure 7: A UML workflow model with required identical components for workflow computational replications highlighted in blue.
@@
This definition asserts that a precisely identical computational workflow produces the same data artifacts at every stage, which is the strictest interpretation of replication similar to the concept of direct replication [54], and is closely related to workflow provenance discussed widely in computational workflow literature [48]. The information required to Replicate a workflow execution computationally is immense. It demonstrates that a specific workflow implementation on specific hardware yields a given set of results, building confidence in workflow implementation but not necessarily the underlying science; it would be challenging for an independent team to Totally Replicate a complex workflow. Recomputation can confirm anomalous or unexpected results before further verification, for instance.
@@
A workflow Totally Replicates another if they Repeat and Reproduce each other and all data artifacts match. Figure 8 contains the generic workflow model with components required for total replication highlighted in blue.
Figure 8: A UML workflow model with required identical components for workflow total replications highlighted in blue.
@@
This definition asserts that a practically identical computational workflow produces the same data artifacts, including interim data, and are identical down to software, but not hardware, details. Total replication is useful to certify a workflow as a ‘gold-standard’ [44] and is similar to the concept of results reproducibility [52]. Despite computational replication being more exacting, we present this tenet last to reflect a balance between exacting replication, feasibility, and the scientific utility in total workflow replication. For example, an integral processing task in radio astronomy requires separating unwanted radio frequency interference from astronomical observations, known as RFI flagging. Not only is the correctness of this task important, but the speed of execution also matters significantly, as this task occurs before any science processing can commence. Many possible methods exist spanning traditional algorithms and machine learning, and a performance claim around a new RFI flagging method could be strengthened by replicating the result independently on different hardware. A total replication permits independent verification of results by a third party and balances scientific and computational replication.
5 Reproducibility Signature Construction Methods
@@
Having provided the theoretical background in the previous section, we outline a practical implementation in this section, allowing us to verify each reproducibility tenant. This includes the specific details behind the workflow signature mechanism and the information captured to build these signatures in DALiuGE, providing a single concrete mapping of the generic workflow model to a workflow management system and the reproducibility tenet definitions.
5.1 Blockchain-Inspired Workflow Signature
@@
Building a dynamic workflow signature is central to our approach to reproducibility testing. Moreover, at SKA scale, these signatures need very efficient construction and comparison routines. A significant challenge of managing computational workflows at this level of scale is the massive parallelism required to handle the volume of data these instruments produce. As such, any workflow operation must handle extreme and dynamic scale in addition to online hardware failures, making robustness and networking efficiency central concerns. To this end, we take inspiration from blockchain technologies, borrowing the Merkle tree cryptographic data structure [57]. Information about each workflow model component, dictated by tenet definitions, is stored in a Merkle tree, the root of which provides a signature for that component. Inserting these signatures into a blockchain-like hash-graph mirroring their connectivity in the original workflow creates a block directed acyclic graph (BlockDAG). The root of a Merkle tree constructed from the leaves of this BlockDAG is a signature for the entire workflow. Similar to how blockchains are cryptographically secure, subtle changes to component provenance information result in drastic changes to the final signature, but structuring this signature makes it efficient to identify where these changes occur. Given two workflow BlockDAGs, one can walk through a topological sort of the nodes to find the first differing signature. The information stored in each block by each workflow component changes for each reproducibility tenet. We collect the hashes of leaf nodes, insert those into a Merkle tree, and take the root as a signature for the entire workflow for a particular tenet.
@@
We use Kahn’s algorithm [58] to traverse a workflow description in a topological order taking $\mathcal{O}(V+E)$ time where $V$ is the number of components (vertices) in the workflow and $E$ is the number of edges between all components. Construction of each component Merkle tree takes $\mathcal{O}(|V|\,log\,|V|)$ time where $|V|$ is the volume of provenance information the component stores. Inserting a component into the BlockDAG takes $\mathcal{O}(d\,log\,d)$ time, where $d$ is the degree of that component. Constructing the entire BlockDAG takes $\mathcal{O}(V(D\,log,D)+E)$ time, where $D$ is the average vertex degree. Construction of the final signature takes $\mathcal{O}(l\,log\,l)$ time, where $l$ is the number of leaves in the workflow graph. When amortized against the time taken to execute a workflow in the first place ($\mathcal{O}(V+E)$), signature construction takes $\mathcal{O}(log\,l)$ time.
5.2 Example Signature Construction
We elaborate on workflow signature construction through an existing simple example [59]. This simple example averages a series of numbers split across multiple tasks, combining the total average in a single process. We present this example using terminology from the generic workflow model but in the phases designated in the life of a DALiuGE workflow, outlined in Section 3.1.2. Figures 9 to 14 provide hash-graphs of a workflow BlockDAG over successive workflow phases under computational replication standards. Information about logical tasks, physical tasks, and logical and physical data artifacts are included to illustrate how this information comes together. Figure 9 presents a logical graph template for a simple averaging workflow and its associated BlockDAG. The original graph structure is maintained, each component generates a hash-value combining its parents’ hash and its own specification.
Figure 9: Example Logical Graph Template (above) and associated hash-graph (below). All components have a one-to-one relation in the BlockDAG. Red boxes represent tasks, blue boxes represent data artifacts, and yellow boxes represent control structures.
A logical graph is a logical template with specified options. Figure 10 presents an example logical graph. The structure has not changed, but more information is incorporated into the signatures, indicated by the ‘fields’ entry, changing the resulting hash values.
Figure 10: Example Logical Graph (above) and associated hash-graph (below). Here, the ‘fields’ entry refers to arbitrary additional information required for a particular tenet. Boxes hold the same meaning as in Figure 9.
Unrolling a logical template into a physical template resolves control structures, yielding a complete description of the computing and storage tasks required. Physical graph templates are directed acyclic graphs. Figure 11 shows a physical graph template where DALiuGE scatters information across two branches.
Figure 11: Example Physical Graph Template (above) and associated hash-graph (below). Control structures have been removed as the workflow has been unrolled. Boxes hold the same meaning as in Figure 9.
Figure 12 shows a physical graph representing a fully specified physical workflow. In this example, physical tasks are spread across multiple machines in a cluster, resulting in principally identical tasks holding different field information (IP address, for example).
Figure 12: Example Physical Graph (above) and associated hash-graph (below). We include the IP address of a particular task’s machine as an example of physical task specific information. Boxes hold the same meaning as in Figure 9.
After runtime, the hash graph contains runtime information such as task completion status and data-artifact hashes. Figure 13 shows this runtime graph.
Figure 13: Example Runtime Graph (above) and associated hash-graph (below). The status field refers to the execution exit code as an example of runtime-specific information. Boxes hold the same meaning as in Figure 9.
Finally, Figure 14 combines the hash-graphs of all workflow components into a final signature. Visualizing more complex workflows does become arduous. However, we aim to show precisely how a signature is created from Logical and physical workflow components.
Figure 14: Example combined hash-graph. The relationships between components throughout their lifecycle, from logical to physical components, are explicit. While the final signatures of each DALiuGE workflow abstraction from LGT to RG are shown separately, in practice, we would hash these signatures together into a single final signature. Boxes hold the same meaning as in Figure 9.
@@ 5.3 Tenet Data
We now elaborate on what specific information each component type in DALiuGE captures for each tenet. The fields we select are ultimately an interpretation of the tenet definitions and, as such, are open to adjustment, especially when considering alternate workflow management systems or native software. The fields selected represent the mapping of the generic workflow model to DALiuGE components, which we have aggregated into ‘Application’ drops, ‘Data’ drops, and ‘Control’ drops for brevity.
Some additional information about DALiuGE is required to understand all fields present. ‘InPorts’ and ‘OutPorts’ represent the DALiuGE-specific abstraction for component dataflow connections. ‘Node’ refers to the logical (integer-valued) or physical (IP address) machine a component will execute on. Status refers to execution status, trace refers to a stack trace of that component’s execution and not all fields are present in all components, as is the case for Filenames in application components, for example. ‘Type’ is an increasingly detailed field as a workflow moves closer to execution. In an LGT, ‘type’ will refer to a generic tool like bash or Python; in a PGT, ‘type’ will refer to a specific toolchain like a Python version. ‘Name’ refers to the exact application name, such as a specific Python script.
@@
Rerunning two workflow executions requires matching their logical workflow components. A DALiuGE mapping implies that logical component details must match. Unsurprisingly, most information captured in this tenet is at the logical layers, only tracking the types and ultimate execution status of each physical component. Table 3 contains a summary of the fields each component type stores in its signature when included in Rerun BlockDAGs.
Table 3: Information stored by DALiuGE components to assert workflow Reruns for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Type
Name
- Status
Data
Type
Storage-Name
Status
Control (misc)
Type
@@
In addition to the information stored for asserting Reruns, Repetition stores additional information about component fields (adjustable parameters). Table 4 summarises the fields present in each component type when included in Repeat BlockDAGs.
Table 4: Information stored to assert workflow repetitions for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Num-CPUs
Fields
Type
Name
- Status
Data
Data-Volume
Fields
Type
Storage-Name
Status
Control (misc) Fields
Type
@@
Recomputation captures significantly more details of component field values and execution environment. This definition and workflow mapping implies that application drops must match exactly. Table 5 summarises the information we capture in testing workflow Recomputation. In addition to the component details, we store the logical partitioning of components to machine resources and an execution trace for each processing component.
Table 5: Information stored to assert workflow Recomputations for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Num-CPUs
Fields
Type
Rank
Name
Node
Island
Node-IP
Island-IP
Status
Trace
Data
Data-Volume
Filenames
Fields
Type
Rank
Storage-Name
Status
Control (misc) Fields
Type
Rank
@@
Reproduction significantly differs, primarily focusing on data components. We store information to assert that data components are configured identically and store the same data summary. The mapping to DALiuGE requires logical data components and data drops to match their specification and content between workflow executions. How components generate their data summary is a subtlety worth noting; in our case, this is a hash of the data contents. In addition to a simple content hash, this content signature could be implemented by calculating statistical properties or other domain-specific characteristics, such as existing hashes in file headers. Table 6 summarises the fields different components store when included in a Reproduction BlockDAG.
Table 6: Information stored to assert workflow Reproductions for different components.
Component-type LGT LG PGT PG RG
Application Type - - - -
Data
Type
Storage-Name
Status
Data-Summary
Control (misc) - -
@@
Replication information is straightforwardly a combination of the Rerun, Recompute, or Repetition information (respectively) and Reproduction fields. Tables 7, 8 and 9 summarise the fields required for scientific, computational, and total replication BlockDAGs, respectively.
Table 7: Information stored to assert workflow Scientific-Replication for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Type
Name
- Status
Data
Type
Storage-Name
Status
Data-Summary
Control (misc)
Type
Table 8: Information stored to assert workflow Computational-Replication for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Num-CPUs
Fields
Type
Rank
Name
Node
Island
Node-IP
Island-IP
Status
Trace
Data
Data-Volume
Filenames
Fields
Type
Rank
Storage-Name
Status
Data-Summary
Control (misc) Fields
Type
Rank
Table 9: Information stored to assert workflow Total-Replication for different components.
Component-type LGT LG PGT PG RG
Application
Type
InPorts
OutPorts
Num-CPUs
Fields
Type
Rank
Name
- Status
Data
Data-Volume
Fields
Type
Storage-Name
Status
Data-Summary
Control (misc) Fields
Type
Rank
Establishing the method used to create a workflow signature and the exact mapping of fields from the generic workflow model to DALiuGE components under different tenet definitions makes the process of workflow signature generation clear.
6 Demonstration
In this section, we fully demonstrate the reproducibility tenets and BlockDAG signature-based testing mechanism. First, we describe a lowpass filter workflow implemented in DALiuGE and a second workflow that tests the effectiveness of a particular lowpass filter. Then, we execute each workflow ten times with different implementations of the lowpass filter element, constructing signatures for each of the reproducibility tenets as described previously. For each tenet, we comment on what the resulting signatures reveal, finding that subtle numerical differences that would otherwise be elusive are obvious when comparing signatures.